OpenStudy (anonymous):

Quick question about trigonometric identities. I have trouble recalling what these mean:

OpenStudy (anonymous):

How do you get \[\sin ^{2}x\] and how do you get \[\cos(2x)\]

OpenStudy (anonymous):

If there is sinx.sinx that will equal \[\sin ^{2}x\]

OpenStudy (anonymous):

or you use the equation \[\cos ^{2}x+\sin ^{2}x=1\] and rearrange to equal \[\sin ^{2}x\]

OpenStudy (anonymous):

You get cos2x from the equation cos(A+B)-cosAcosB-sinAsinB And you let B=A Which gives cos2A=cosAcosA-sinAsinA

OpenStudy (anonymous):

Thanks @WilliamF

OpenStudy (anonymous):

your welcome